# Scaling regimes for unsteady diffusion across particle-stabilized fluid interfaces

### v2

@@ Abstract
In this Letter, we reconcile the apparent contradictions in the effect of particle-laden interfaces on mass transfer, using an unsteady diffusion model that accounts for the particles through a reduction in available area. The resulting diffusive transport closely resembles that across a homogeneous layer with reduced effective diffusivity. Through this analogy, we demonstrate how diffusion can appear unaffected by a particle layer or barrier, even when local diffusivity or interfacial area is significantly reduced. Our approach further unifies transport phenomena across diverse systems, including multicomponent droplet evaporation and vapor dissolution, as shown in Fig. 1.
@@
Systems showing hindered transport due to the interfacial layer (filled symbols) predominantly lie in the region where $t_{L}>R^{2}/D_{0}$ (shaded area), while cases where the layer has no effect fall well outside this region. This parameter space overview highlights that effective hindered transport requires interfaces that are both sufficiently thick (typically $a/R\gtrsim 0.01$) and very densely covered (with most data points for $\phi_{0}\gtrsim 0.99$). For particle monolayers, this demands exceptionally high surface coverage, well above the dense circle packing limit $\phi\gg 0.91$. In the experiments reported in Fig. 4, this is achieved, for instance, by thermal sintering [18] or polymer deposition [15]. For armored bubbles, the literature typically reports only the late stages of the system, after dissolution arrest [11, 34] or buckling [35, 36], providing insufficient data to include them in our analysis in Fig. 4.
The results from our transient approach are consistent with Berg’s steady-state solutions: both frameworks predict that diffusive transport is significantly hindered for very high surface coverage [22]. Yet, a transient approach is especially important for well-covered interfaces, where the timescale $t_{L}$, marking the onset of bulk-dominated diffusion, can exceed the layer diffusion timescale $a^{2}/D_{L}$ by multiple orders of magnitude. As a result, transients may persist throughout all experimentally relevant time scales, rendering steady-state assumptions invalid. This time-resolved perspective qualitatively distinguishes between experiments with and without long-term transport hindrance. While such a simplified model cannot capture the full transient dynamics of each system, the effect of one dominant mechanism governing mass transfer across a porous layer seems to be sufficient to resolve the contrasting observations in the literature. As a future direction, more accurate models including specific physico-chemical properties of emulsions, evaporating drops, or dissolving bubbles will enable more quantitative descriptions of mass transfer across complex interfaces. More broadly, the model is generally applicable to other complex interfaces ubiquitous in nature and technology.

### v3

@@ Abstract
In this Letter, we reconcile the apparent contradictions in the effect of particle-laden interfaces on mass transfer, using an unsteady diffusion model that accounts for the particles through a reduction in available area. The resulting diffusive transport closely resembles that across a homogeneous layer with reduced effective diffusivity. Through this analogy, we demonstrate how diffusion can appear unaffected by a particle layer or barrier, even when local diffusivity or interfacial area is significantly reduced. Even though our simplified model does not account for detailed physicochemical processes, it captures a dominant mechanism that further unifies transport phenomena across diverse systems, including multicomponent droplet evaporation and vapor dissolution, as shown in Fig. 1.
@@
Systems showing hindered transport due to the interfacial layer (filled symbols) predominantly lie in the region where $t_{L}>R^{2}/D_{0}$ (shaded area), while cases where the layer has no effect fall well outside this region. Note that the threshold remains inconclusive for some cases, particularly the gray and white symbols within the shaded areas. This may reflect the uncertainties in estimating layer thickness and surface coverage from specific experiments, and could also suggest that additional physical effects are relevant in those cases. Nonetheless, the parameter space overview highlights that effective hindered transport requires interfaces that are both sufficiently thick (typically $a/R\gtrsim 0.01$) and very densely covered (with most data points for $\phi_{0}\gtrsim 0.99$). For particle monolayers, this demands exceptionally high surface coverage, well above the dense circle packing limit $\phi\gg 0.91$. In the experiments reported in Fig. 4, this is achieved, for instance, by thermal sintering [18] or polymer deposition [15]. For armored bubbles, the literature typically reports only the late stages of the system, after dissolution arrest [11, 34] or buckling [35, 36], providing insufficient data to include them in our analysis in Fig. 4.
The results from our transient approach are consistent with Berg’s steady-state solutions: both frameworks predict that diffusive transport is significantly hindered for very high surface coverage [22]. Yet, a transient approach is especially important for well-covered interfaces, where the timescale $t_{L}$, marking the onset of bulk-dominated diffusion, can exceed the layer diffusion timescale $a^{2}/D_{L}$ by multiple orders of magnitude. As a result, transients may persist throughout all experimentally relevant time scales, rendering steady-state assumptions invalid. This time-resolved perspective qualitatively distinguishes between experiments with and without long-term transport hindrance. While such a simplified model cannot capture the full transient dynamics of each system, the results highlight the major effect of diffusion to explain the contrasting observations in the literature. As a future direction, more accurate models including specific physico-chemical properties of emulsions, evaporating drops, or dissolving bubbles will enable more quantitative descriptions of mass transfer across complex interfaces. More broadly, the model is generally applicable to other complex interfaces ubiquitous in nature and technology.
@@ Acknowledgements.
The data that support the findings of this article are openly available [37].